OpenStudy (anonymous):

what the point slope equation of the line with 3/2. that goes through the point (-3,5) ?

OpenStudy (anonymous):

y-y1 = m(x-x1) m= slope

OpenStudy (hba):

\[y-y_1=m(x-x_1)\] where, m=3/2 (x1,y1)=(-3,5)

OpenStudy (anonymous):

y + 5 = -3/2(x-3) ?

OpenStudy (hba):

m=3/2 not -3/2

OpenStudy (hba):

So, y-5=3/2(x+3)

OpenStudy (hba):

Now you can take 2 on that side 2(y-5)=3(x+3) now solve